# Reasoning Under Uncertainty: Exploring Probabilistic Reasoning Capabilities of LLMs

### v1

@@ 5.4.1 Label Sensitivity Results
To investigate the impact of label choice on model performance, a targeted set of experiments is conducted on three models: the two top-performing models, Llama3.3-70B and GPT-4.1-mini, and a smaller model, Qwen2.5-7B. All three are evaluated on the joint-mode task, which they handle reliably under standard conditions, using joint space size of $|\mathcal{L}|=54$ and ten distinct label categories. These categories span a broad range of semantics, including computer science terms, human names, US states, countries, fruits, universities, majors, car brands, neurology terms, and food. By systematically varying the labels used to represent the outcomes of the random variables, the goal is to assess how surface-level changes in label representation influence model behavior.
Figure 3 (right) presents the accuracy of each model across the different label categories. GPT-4.1-mini shows strong consistency, with minimal variation in accuracy across categories, indicating a high level of robustness to changes in label sets. Llama3.3-70B performs well overall but exhibits moderate sensitivity, with performance varying more noticeably across label sets. In contrast, Qwen2.5-7B is highly sensitive to label changes, suffering substantial drops in accuracy depending on the category. These findings suggest that while larger models or those distilled from them can generalize well across prompt variations, smaller models are significantly vulnerable to the way information is presented, even when the underlying task remains unchanged.
@@ 5.4.2 Samples-in-Context Results
Next, we investigate how model performance is affected when raw observed samples are provided in the prompt instead of their empirical frequencies. This setting is evaluated using the two best-performing models, Llama3.3-70B and GPT-4.1-mini, as smaller models consistently fail under these conditions. We repeat the joint-mode task across different joint space sizes, but rather than supplying frequency counts for each outcome, we provide the raw samples directly as mentioned. This forces the models to perform implicit counting—first identifying the frequency of each outcome from the raw data, and then determining the mode based on that.
The results reveal a sharp decline in performance: both models, which had previously achieved near-perfect accuracy with explicit frequencies, struggle significantly in this setting. As the joint space size $|\mathcal{L}|$ increases and the input grows longer, their ability to track counts and identify the correct outcome diminishes rapidly. This performance drop is illustrated in Figure 3 (left), which shows a near-linear decline in accuracy with increasing joint size. The findings point to a core limitation: the models’ inability to reliably perform counting over long contexts. Among the two, GPT-4.1-mini consistently outperforms Llama3.3-70B, indicating greater robustness to the added complexity and input length. Further evidence from Appendix F shows that even when the joint space size is fixed at $|\mathcal{L}|=12$, increasing the number of observed samples $K$ results in additional accuracy degradation. This suggests that the challenge is not only tied to task complexity but also to the models’ limited capacity for reasoning over long sequences.
To explore potential mitigation strategies, we conducted an additional experiment with GPT-4.1-mini where the model was given access to a code interpreter. In this setup, the model was encouraged to write Python code to count frequencies before solving the task. This hybrid approach substantially improved performance, nearly matching the results from the frequency-provided setting. The detailed results of these experiments are presented in Appendix G.
@@ 6 Conclusion
This work provides the first comprehensive evaluation of large language models’ abilities to reason over discrete probability distributions. By examining a range of structured tasks including mode identification, maximum likelihood estimation, and sample generation, we observe clear differences in capability between models. Larger models or those distilled from them generally perform well across tasks, but their strengths are especially evident in sample generation, where the produced samples align more closely with the target distribution than those of a true random sampler. However, we find that performance consistently degrades when tasks are framed as conditional rather than joint distributions, revealing an ongoing challenge in conditional reasoning. Moreover, models remain sensitive to superficial changes in outcome labels and consistently fail when required to infer probabilities from raw samples instead of provided frequencies. Together, these findings highlight both the significant progress of modern LLMs in probabilistic reasoning and the fundamental challenges that must be addressed to achieve reliable reasoning under uncertainty.
@@ Acknowledgements
This section includes complete prompt examples for each task, the exact frequency distributions used for various joint space sizes, the prompt template used to evaluate outputs with the LLM judge and supplemental results for samples-in-Context settings.
@@ A.1 Mode Identification
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are:
- (A, C, E): 5
- (A, C, F): 7
- (A, C, G): 3
- (A, D, E): 10
- (A, D, F): 2
- (A, D, G): 3
- (B, C, E): 6
- (B, C, F): 8
- (B, C, G): 2
- (B, D, E): 3
- (B, D, F): 3
- (B, D, G): 8
Task: Identify the mode (most probable outcome) of the joint distribution P(X, Y, Z).
Instructions:
1. Do not write any code or pseudocode.
2. Strictly follow the output format below.
3. You may explain your reasoning, but the final answer should be explicitly summarized at the end. You get negative penalty for not following the output format.
Output Format that you should strictly follow:
Mode = (, , )
@@
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are:
- (A, C, E): 5
- (A, C, F): 7
- (A, C, G): 3
- (A, D, E): 10
- (A, D, F): 2
- (A, D, G): 3
- (B, C, E): 6
- (B, C, F): 8
- (B, C, G): 2
- (B, D, E): 3
- (B, D, F): 3
- (B, D, G): 8
Task: Identify the mode (most probable outcome) of the conditional distribution P(Z | X = A).
Instructions:
1. Do not write any code or pseudocode.
2. Strictly follow the output format below.
3. You may explain your reasoning, but the final answer should be explicitly summarized at the end. You get negative penalty for not following the output format.
Output Format that you should strictly follow:
Most probable value of Z given X=A is
@@ A.2 Maximum Likelihood Estimation
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are:
- (A, C, E): 5
- (A, C, F): 7
- (A, C, G): 3
- (A, D, E): 10
- (A, D, F): 2
- (A, D, G): 3
- (B, C, E): 6
- (B, C, F): 8
- (B, C, G): 2
- (B, D, E): 3
- (B, D, F): 3
- (B, D, G): 8
Task: Predict the maximum likelihood estimation (MLE) of the joint probability distribution P(X, Y, Z) based on these 60 samples.
Instructions:
1. Think step by step and solve this mathematically using probability theory - do not write any code or pseudocodes as you get negative penalty for that.
2. Clearly state the final estimated probabilities for each (X, Y, Z) outcome. Output probabilities should be expressed as float numbers with up to four decimal points.
3. You may explain your reasoning, but the final answer should be explicitly summarized at the end. only your final answer will be graded.
Output Format:
- Final probabilities as float numbers should be listed as:
P(A, C, E) = [value]
P(A, C, F) = [value]
P(A, C, G) = [value]
P(A, D, E) = [value]
P(B, D, F) = [value]
P(B, D, G) = [value]
@@
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are:
- (A, C, E): 5
- (A, C, F): 7
- (A, C, G): 3
- (A, D, E): 10
- (A, D, F): 2
- (A, D, G): 3
- (B, C, E): 6
- (B, C, F): 8
- (B, C, G): 2
- (B, D, E): 3
- (B, D, F): 3
- (B, D, G): 8
Task: Predict the maximum likelihood estimation (MLE) of the conditional distribution P(X | Z = E).
Instructions:
1. Think step by step and solve this mathematically using probability theory - do not write any code or pseudocodes as you get negative penalty for that.
2. Clearly state the final estimated probabilities as float numbers with up to four decimal points.
3. You may explain your reasoning, but the final answer should be explicitly summarized at the end.
Output Format:
- Final probabilities as float numbers should be listed as:
P(X=A | Z=E) = [value]
P(X=B | Z=E) = [value]
@@ A.3 Sample Generation
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are:
- (A, C, E): 5
- (A, C, F): 7
- (A, C, G): 3
- (A, D, E): 10
- (A, D, F): 2
- (A, D, G): 3
- (B, C, E): 6
- (B, C, F): 8
- (B, C, G): 2
- (B, D, E): 3
- (B, D, F): 3
- (B, D, G): 8
Task: Generate EXACTLY 84 random samples from the joint distribution P(X, Y, Z).
Instructions:
1. OUTPUT MUST BEGIN with ### Output on a new line.
2. List EXACTLY 84 samples numbered 1 to 84.
3. Do not write any code or pseudocode.
4. Strictly follow the output format below.
5. You will be penalized for writing codes or not following the output format.
Output Format that you should strictly follow:
### Output
1. (X, Y, Z)
2. (X, Y, Z)
84. (X, Y, Z)
Where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G.
@@
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are:
- (A, C, E): 5
- (A, C, F): 7
- (A, C, G): 3
- (A, D, E): 10
- (A, D, F): 2
- (A, D, G): 3
- (B, C, E): 6
- (B, C, F): 8
- (B, C, G): 2
- (B, D, E): 3
- (B, D, F): 3
- (B, D, G): 8
Task: Generate EXACTLY 84 random samples from the conditional distribution P(Z | X=A).
Instructions:
1. OUTPUT MUST BEGIN with ### Output on a new line.
2. List EXACTLY 84 samples numbered 1 to 84.
3. Do not write any code or pseudocodes.
4. Strictly follow the output format below.
5. You will be penalized for writing codes or not following the output format.
Output Format that you should strictly follow:
### Output
1. Z
2. Z
84. Z
Where Z can take outcomes from E, G, F.
@@ Appendix C LLM Judge Prompt
We used the LLM judge only for the mode identification task, where model responses were often difficult to evaluate using rule-based methods. Many responses implied an answer but didn’t state it clearly or follow a consistent format, making extraction with regular expressions unreliable. For the other two tasks—MLE and sampling—models typically followed the expected output format, or when they didn’t, their responses clearly lacked an answer, making them easy to discard.
@@
And here’s the template of the user prompt that is given to the judge model:
Given the following question, expected answer, and response, judge whether the response is correct. Clearly state your judgment as "Judgment: Correct." or "Judgment: Incorrect.".
Question: [question]
Expected Answer: [expected]
Response to Evaluate: [response]
Judgment (Correct/Incorrect):
Appendix D Independence of Generated Samples

### v2

@@ 5.4.1 Label Sensitivity Results
To investigate the impact of label choice on model performance, a targeted set of experiments is conducted on three models: the two top-performing models, Llama3.3-70B and GPT-4.1-mini, and a smaller model, Qwen2.5-7B. All three are evaluated on the joint-mode task, which they handle reliably under standard conditions, using joint space size of $|\mathcal{L}|=54$ and ten distinct label categories. These categories span a broad range of semantics, including computer science terms, human names, US states, countries, fruits, universities, majors, car brands, neurology terms, and food. By systematically varying the labels used to represent the outcomes of the random variables, we aimed to assess how surface-level changes in label representation influence model behavior.
Figure 3 (right) presents the accuracy of each model across the different label categories. GPT-4.1-mini shows remarkable consistency, with minimal variation across labels, indicating strong robustness. Llama3.3-70B performs well overall but exhibits moderate sensitivity, with performance varying more noticeably across label sets. In contrast, Qwen2.5-7B is highly sensitive to label changes, suffering sharp drops in accuracy depending on the category. These findings suggest that while larger models or those distilled from them can generalize well across prompt variations, smaller models are significantly vulnerable to the way information is presented, even when the underlying task remains unchanged.
@@ 5.4.2 Samples-in-Context Results
Next, we investigate how model performance is affected when raw observed samples are provided in the prompt instead of their empirical frequencies. This setting is evaluated using the two best-performing models, Llama3.3-70B and GPT-4.1-mini, as smaller models consistently fail under these conditions. We repeat the joint-mode task across different joint space sizes, but rather than supplying frequency counts for each outcome, we provide the raw samples directly as mentioned. This forces the models to perform implicit counting, first identifying the frequency of each outcome from the raw data, and then determining the mode based on that.
The empirical evaluations reveal a sharp decline in performance. Both models, which had previously achieved near-perfect accuracy with explicit frequencies, struggle significantly in this setting. As the joint space size $|\mathcal{L}|$ increases and the input grows longer, their ability to track counts and identify the correct outcome diminishes rapidly. Figure 3 (left) illustrates this trend, showing an almost linear decrease in accuracy with increasing joint size. Among the two, GPT-4.1-mini consistently outperforms Llama3.3-70B, indicating greater robustness to the added complexity and input length. Further evidence from Appendix G shows that even when the joint space size is fixed at $|\mathcal{L}|=12$, increasing the number of observed samples $K$ results in further degradation of accuracy. This suggests that the challenge is not only tied to task complexity but also to the models’ limited capacity for reasoning over long sequences.
To explore potential mitigation strategies, we conducted an additional experiment with GPT-4.1-mini where the model was given access to a code interpreter. In this setup, the model was encouraged to write Python code to count frequencies before solving the task. This hybrid approach substantially improved performance, nearly matching the results from the frequency-provided setting. The detailed results of these experiments are reported in Appendix H.
@@ 6 Conclusion
This work presents the first comprehensive evaluation of large language models’ abilities to understand discrete probability distributions. Through a set of structured tasks, mode identification, maximum likelihood estimation, and sample generation, we reveal clear differences in performance across models. Larger models, or those distilled from them, generally perform well, with sample generation emerging as a surprising strength, where in some cases, generated samples align with the target distribution even more closely than those from a true random sampler. Yet, notable weaknesses remain. Performance consistently declines when tasks involve conditional distributions rather than joint distributions, reflecting the difficulty of conditional reasoning. Models also show sensitivity to superficial changes in the outcomes of the random variables and struggle when asked to reason over long contexts or to infer probabilities directly from raw samples. Taken together, these findings suggest that while pretraining knowledge enables LLMs to handle probabilistic tasks, their performance declines as the complexity of the problems grows. In such cases, providing additional support, such as in-context examples or external tools like a code interpreter, can help boost the performance. Addressing these challenges is a key step toward developing LLMs capable of robust reasoning under uncertainty.
@@ Acknowledgements
This section provides additional details, prompts, and results that supplement the main text. Section A presents complete prompt examples for each task, while Section B lists the exact frequency distributions used for different joint space sizes. Section C describes the LLM judge prompt template employed to evaluate outputs. Section D explores one-shot prompting, including examples and results, and Section E examines whether samples generated by LLMs are independent. Section F reports evaluations on a real-world dataset to assess generalizability, Section G provides supplemental results on the effect of sample count in context and Section H details experiments where GPT-4.1-mini was given access to a code interpreter.
@@ A.1 Mode Identification
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are: - (A, C, E): 5 - (A, C, F): 7 - (A, C, G): 3 - (A, D, E): 10 - (A, D, F): 2 - (A, D, G): 3 - (B, C, E): 6 - (B, C, F): 8 - (B, C, G): 2 - (B, D, E): 3 - (B, D, F): 3 - (B, D, G): 8 Task: Identify the mode (most probable outcome) of the joint distribution P(X, Y, Z). Instructions: 1. Do not write any code or pseudocode. 2. Strictly follow the output format below. 3. You may explain your reasoning, but the final answer should be explicitly summarized at the end. You get negative penalty for not following the output format. Output Format that you should strictly follow: Mode = (, , )
@@
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are: - (A, C, E): 5 - (A, C, F): 7 - (A, C, G): 3 - (A, D, E): 10 - (A, D, F): 2 - (A, D, G): 3 - (B, C, E): 6 - (B, C, F): 8 - (B, C, G): 2 - (B, D, E): 3 - (B, D, F): 3 - (B, D, G): 8 Task: Identify the mode (most probable outcome) of the conditional distribution P(Z | X = A). Instructions: 1. Do not write any code or pseudocode. 2. Strictly follow the output format below. 3. You may explain your reasoning, but the final answer should be explicitly summarized at the end. You get negative penalty for not following the output format. Output Format that you should strictly follow: Most probable value of Z given X=A is
@@ A.2 Maximum Likelihood Estimation
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are: - (A, C, E): 5 - (A, C, F): 7 - (A, C, G): 3 - (A, D, E): 10 - (A, D, F): 2 - (A, D, G): 3 - (B, C, E): 6 - (B, C, F): 8 - (B, C, G): 2 - (B, D, E): 3 - (B, D, F): 3 - (B, D, G): 8 Task: Predict the maximum likelihood estimation (MLE) of the joint probability distribution P(X, Y, Z) based on these 60 samples. Instructions: 1. Think step by step and solve this mathematically using probability theory - do not write any code or pseudocodes as you get negative penalty for that. 2. Clearly state the final estimated probabilities for each (X, Y, Z) outcome. Output probabilities should be expressed as float numbers with up to four decimal points. 3. You may explain your reasoning, but the final answer should be explicitly summarized at the end. only your final answer will be graded. Output Format: - Final probabilities as float numbers should be listed as: P(A, C, E) = [value] P(A, C, F) = [value] P(A, C, G) = [value] P(A, D, E) = [value] P(B, D, F) = [value] P(B, D, G) = [value]
@@
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are: - (A, C, E): 5 - (A, C, F): 7 - (A, C, G): 3 - (A, D, E): 10 - (A, D, F): 2 - (A, D, G): 3 - (B, C, E): 6 - (B, C, F): 8 - (B, C, G): 2 - (B, D, E): 3 - (B, D, F): 3 - (B, D, G): 8 Task: Predict the maximum likelihood estimation (MLE) of the conditional distribution P(X | Z = E). Instructions: 1. Think step by step and solve this mathematically using probability theory - do not write any code or pseudocodes as you get negative penalty for that. 2. Clearly state the final estimated probabilities as float numbers with up to four decimal points. 3. You may explain your reasoning, but the final answer should be explicitly summarized at the end. Output Format: - Final probabilities as float numbers should be listed as: P(X=A | Z=E) = [value] P(X=B | Z=E) = [value]
@@ A.3 Sample Generation
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are: - (A, C, E): 5 - (A, C, F): 7 - (A, C, G): 3 - (A, D, E): 10 - (A, D, F): 2 - (A, D, G): 3 - (B, C, E): 6 - (B, C, F): 8 - (B, C, G): 2 - (B, D, E): 3 - (B, D, F): 3 - (B, D, G): 8 Task: Generate EXACTLY 84 random samples from the joint distribution P(X, Y, Z). Instructions: 1. OUTPUT MUST BEGIN with ### Output on a new line. 2. List EXACTLY 84 samples numbered 1 to 84. 3. Do not write any code or pseudocode. 4. Strictly follow the output format below. 5. You will be penalized for writing codes or not following the output format. Output Format that you should strictly follow: ### Output 1. (X, Y, Z) 2. (X, Y, Z) 84. (X, Y, Z) Where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G.
@@
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are: - (A, C, E): 5 - (A, C, F): 7 - (A, C, G): 3 - (A, D, E): 10 - (A, D, F): 2 - (A, D, G): 3 - (B, C, E): 6 - (B, C, F): 8 - (B, C, G): 2 - (B, D, E): 3 - (B, D, F): 3 - (B, D, G): 8 Task: Generate EXACTLY 84 random samples from the conditional distribution P(Z | X=A). Instructions: 1. OUTPUT MUST BEGIN with ### Output on a new line. 2. List EXACTLY 84 samples numbered 1 to 84. 3. Do not write any code or pseudocodes. 4. Strictly follow the output format below. 5. You will be penalized for writing codes or not following the output format. Output Format that you should strictly follow: ### Output 1. Z 2. Z 84. Z Where Z can take outcomes from E, G, F.
@@ Appendix C LLM Judge Prompt
We used the LLM judge only for the mode identification task, where model responses were often difficult to evaluate using rule-based methods. Many responses implied an answer but did not state it clearly or follow a consistent format, making extraction with regular expressions unreliable. For the other two tasks—MLE and sampling—models typically followed the expected output format, or when they did not, their responses clearly lacked an answer, making them easy to discard.
@@
The template of the user prompt that is provided to the judge model is given below:
Given the following question, expected answer, and response, judge whether the response is correct. Clearly state your judgment as "Judgment: Correct." or "Judgment: Incorrect.". Question: [question] Expected Answer: [expected] Response to Evaluate: [response] Judgment (Correct/Incorrect):
Appendix D One-shot Prompting
Table 3: Performance comparison of different models in zero-shot vs one-shot settings across four tasks with $|\mathcal{L}|=54$. Mode identification tasks are evaluated using accuracy, while MLE tasks are evaluated using total variation distance (TVD).
joint-mode cond-mode joint-MLE cond-MLE
Model zero-shot one-shot zero-shot one-shot zero-shot one-shot zero-shot one-shot
Llama3.1-8B 0.21 0.88 0.62 0.93 0.02 0 0.198 0.082
Qwen2.5-7B 0.65 0.98 0.60 0.87 0.006 6e-05 0.177 0.041
DeepSeek-R1Distill-Qwen-7B 0.86 0.9 0.83 0.95 0.02 0.001 0.100 0.09
Llama3.3-70B 0.97 1.0 0.95 1.0 4e-04 0 0.084 0.004
GPT-4o-mini 0.80 1.0 0.88 1.0 1e-04 0 0.084 0.003
GPT-4.1-mini 0.96 1.0 1.0 1.0 1e-04 0 0.016 3e-04
As illustrated in Tables 1 and  2, models tend to struggle on the more complex tasks when relying solely on their pretraining knowledge. To address this, we designed a set of experiments in which each prompt was augmented with a single, simpler example of the task. Specifically, we included an example with $|\mathcal{L}|=12$ along with its step-by-step solution, followed by the target, more challenging task with $|\mathcal{L}|=54$, and asked the model to solve it.
As shown in Table 3, this one-shot prompting strategy substantially improved the models’ accuracy, highlighting the impact of providing minimal guidance on solving the harder problems. Interestingly, we observed that the performance of smaller models with just one in-context example closely matched the zero-shot performance of much larger models, underscoring the strength of in-context learning as a way to compensate for smaller model size.
An example of the one-shot prompt used in our experiments is provided below.
Task Definition: Consider the joint probability distribution of a set of discrete random variables. You will be provided with the frequencies of each outcome of the joint distribution in a set of independently drawn samples. Your task is to identify the mode (the most probable outcome) of a specific conditional distribution. — EXAMPLE with Solution— Consider discrete random variables X, Y, Z with joint distribution P(X, Y, Z) where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from the joint distribution P(X, Y, Z), the observed frequencies are: - (A, C, E): 5 - (A, C, F): 7 - (A, C, G): 3 - (A, D, E): 10 - (A, D, F): 2 - (A, D, G): 3 - (B, C, E): 6 - (B, C, F): 8 - (B, C, G): 2 - (B, D, E): 3 - (B, D, F): 3 - (B, D, G): 8 Example Task (solved): Identify the mode (most probable outcome) of the conditional distribution P(Z | X = A). Step-by-step solution: 1. Extract rows where X = A: (A,C,E):5, (A,C,F):7, (A,C,G):3, (A,D,E):10, (A,D,F):2, (A,D,G):3. 2. Aggregate counts by Z among those rows: - Count(Z = E | X = A) = 5 + 10 = 15. - Count(Z = F | X = A) = 7 + 2 = 9. - Count(Z = G | X = A) = 3 + 3 = 6. 3. Total samples with X = A = 15 + 9 + 6 = 30. 4. Compute empirical conditional probabilities. - P(Z = E | X = A) = 15 / 30 = 0.5 - P(Z = F | X = A) = 9 / 30 = 0.3 - P(Z = G | X = A) = 6 / 30 = 0.2 5. Identify the mode: Z = E (highest probability 0.5). Final Answer: Most probable value of Z given X=A is E — END EXAMPLE — Main Task: Consider 4 discrete random variables X, Y, Z, T, where X can take outcomes from A, B, C, Y can take outcomes from D, E, Z can take outcomes from F, G, H, T can take outcomes from I, J, K. In 270 independent samples drawn from their joint distribution P(X, Y, Z, T), the observed frequencies are: - (A, D, F, I): 6 - (A, D, F, J): 3 - (C, E, H, I): 4 - (C, E, H, J): 2 - (C, E, H, K): 1 Main Task (to solve): Identify the mode (most probable outcome) of the conditional distribution P(X | Y = E). Instructions: 1. Do not write any code or pseudocodes. 2. Strictly follow the output format below. 3. You may explain your reasoning, but the final answer should be explicitly summarized at the end. You get negative penalty for not following the output format. Output Format that you should strictly follow: Most probable value of X given Y=E is
Appendix E Independence of Generated Samples